OpenStudy (anonymous):

if i have a 91.7% grade and get a 0/80 on a test, what will my grade be then

OpenStudy (carterjackson):

this all depends on the weight score for tests in the class (ex. 40% tests 20% quizzes 40% homework)

OpenStudy (anonymous):

tests are worth more i think

OpenStudy (jtug6):

You need to know the averages and then multiply those by the weighted percentages of your exams/quizzes/hw ect. So sum up the quizzes/hw since those wont change then take your current score and add 0 then divide it by 80. So if you have a 90/100 for tests so far then (90+0)/(100+80) = 90/180 = 1/2 so then multiply that by the weight of your test and then sum that to your quiz/hw and that'll give you your grade.

OpenStudy (jtug6):

As an example: 40% for tests 30% for quizzes 30% for hw If i'm a 90/100 for tests then (90/100) * 40 = 36 If i'm a 85/100 for quizzes then (85/100) * 30 = 25.5 If i'm 100/100 for homework then (100/100) * 30 = 30 Now sum them all up: 36 + 25.5 + 30 = 91.5

OpenStudy (chenclan):

Depends on how much assignment you have done like if you did 8000 assignments it wont affect you much not even 1%